OpenStudy (anonymous):

Find f'(x) and f'(c). MEDAL WILL BE REWARDED!!! Function: f(x)=(sinx)/x Value of c: pi/6

myininaya (myininaya):

Do you know the quotient rule?

OpenStudy (anonymous):

If I can get the equation in line then yes

myininaya (myininaya):

in line?

OpenStudy (anonymous):

In final form

myininaya (myininaya):

I don't understand what you mean. This is a quotient because there is a division sign. I'm asking do you know a particular rule. The quotient rule is the one I'm asking about. If you have (f/g)' then by quotient rule this is [f'*g-g'*f]/g^2

OpenStudy (anonymous):

So would I get (sinx-cosx)/x^2

myininaya (myininaya):

\[(\frac{f}{g})'=\frac{f' \cdot g- g' \cdot f}{g^2}\] no... Like f is the top and your top here is sin(x) and g is the bottom and your bottom is x You are going to go in that new quotient I provided for you and plug those in. you also need to plug in for f' and g'.

OpenStudy (anonymous):

okay, so sinx(x)-x^2(sinx)/x^2

myininaya (myininaya):

\[(\frac{\sin(x)}{x})'=\frac{(\sin(x))' \cdot x -(x)' \cdot \sin(x)}{x^2} \] like this....see i replaced my g's in that formula with x and I replaced my f's with sin(x)

myininaya (myininaya):

Those little prime things tells us we need to find the derivative though .

OpenStudy (anonymous):

oh okay give me one sec ill decipher this...

OpenStudy (anonymous):

Im stuck, if I start from scrath I get sin(X)

myininaya (myininaya):

Ok... Do you understand how I get replaced the f items with sin(x) and the g items with x?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

Ok so the only thing to do is just find the derivative of sin(x) to replace (sin(x))' and just to find the derivative x to replace (x)' in: \[(\frac{\sin(x)}{x})'=\frac{(\sin(x))' \cdot x -(x)' \sin(x)}{x^2} \]

myininaya (myininaya):

what is the derivative of sin(x)?

OpenStudy (anonymous):

cos(x)

myininaya (myininaya):

so now we \[(\frac{\sin(x)}{x})'=\frac{\cos(x) \cdot x - (x)' \sin(x)}{x^2} \]

myininaya (myininaya):

what is the derivative of x?

OpenStudy (anonymous):

1

myininaya (myininaya):

so this means we have \[(\frac{\sin(x)}{x})'=\frac{x \cos(x) - 1 \sin(x)}{x^2} =\frac{xcos(x)-\sin(x)}{x^2} \] since 1sin(x)=sin(x)

myininaya (myininaya):

now your last direction says to evaluate this at pi/6

myininaya (myininaya):

Before we move on, are there any questions?

OpenStudy (anonymous):

no I understand this now! Lets keep moving! :)

myininaya (myininaya):

\[(\frac{\sin(x)}{x})'|_{x=\frac{\pi}{6}}=\frac{\frac{\pi}{6} \cos(\frac{\pi}{6})-\sin(\frac{\pi}{6})}{{(\frac{\pi}6})^2} \]

myininaya (myininaya):

Do you know how to use the unit circle to evaluate those trig parts?

OpenStudy (anonymous):

No, that's my biggest problem!

myininaya (myininaya):

Now you are looking on that circle for when pi/6 occurs.

myininaya (myininaya):

Do you see pi/6?

OpenStudy (anonymous):

Okay so what exactly am I looking for.....yes I do! at the first 30?

myininaya (myininaya):

Do you see the ordered pair when we have pi/6 radians?

OpenStudy (anonymous):

yes, is that my final answer?

myininaya (myininaya):

That first value in there or the first coordinate is your cos(pi/6) value. That second value in there or the second coordinate is your sin(pi/6) value.

OpenStudy (anonymous):

so plug them in to the equation?

myininaya (myininaya):

first coordinate meaning x-coordinate second coordinate meaning y-coordinate

myininaya (myininaya):

yep

myininaya (myininaya):

tell me what you have when you plug them in

OpenStudy (anonymous):

okay hold on with me....give me one sec

OpenStudy (anonymous):

wait what will the equation look like?

myininaya (myininaya):

Tell me what cos(pi/6) is looking on that unit circle?

OpenStudy (anonymous):

radical3/2, 1/2

OpenStudy (anonymous):

and 1/2 is our sin(pi/6) value?

myininaya (myininaya):

both of those aren't cos(pi/6)

myininaya (myininaya):

ok

myininaya (myininaya):

so cos(pi/6) is sqrt(3)/2 and sin(pi/6) is 1/2 now I'm going to go to my formula and replace those trig parts with these lovely fractions...

OpenStudy (anonymous):

so would it be radical3/2(1/2) - (1/2)/radical3/2^2?

myininaya (myininaya):

\[(\frac{\sin(x)}{x})'|_{x=\frac{\pi}{6}}=\frac{\frac{\pi}{6} \cos(\frac{\pi}{6})-\sin(\frac{\pi}{6})}{{(\frac{\pi}6})^2} \] \[=\frac{\frac{\pi}{6} \cdot \frac{\sqrt{3}}{2}-\frac{1}{2}}{(\frac{\pi}{6})^2} \]

myininaya (myininaya):

I replaced cos(pi/6) with sqrt(3)/2 and I replaced sin(pi/6) with 1/2 and did nothing else

OpenStudy (anonymous):

okay, I see now, I had my equation written down wrong, so now I can just solve?

myininaya (myininaya):

we don't call it solving we call it evaluating/simplifying an expression for a certain value

myininaya (myininaya):

but yes this can be simplified

OpenStudy (anonymous):

okay so can I simplify?

myininaya (myininaya):

\[\frac{ \frac{\sqrt{3} \pi}{12}-\frac{1}{2}}{\frac{\pi^2}{36}}\] I would start by clearing the compound fractions... Meaning let's get rid of the mini-fractions in our big fraction.

OpenStudy (anonymous):

I get 3(pi radical3 -6)/6

OpenStudy (anonymous):

over pi I mean

myininaya (myininaya):

What can we multiply on top and bottom for our big (main) fraction so that we get rid of the the mini-fractions ? Look at my last fraction I put up.

OpenStudy (anonymous):

you can get rid of the pi's or the GCF

myininaya (myininaya):

Look our mini-fractions denominators 12,2, and 36. The lcm of 12, 2, and 36 sometimes we denote this as lcm(12,2,36) is 36 correct?

OpenStudy (anonymous):

yes

myininaya (myininaya):

lcm meaning least common multiple

myininaya (myininaya):

so multiply top and bottom of my main fraction by 36

myininaya (myininaya):

\[\frac{36}{36} \cdot \frac{ \frac{\sqrt{3} \pi }{12} -\frac{1}{2}}{\frac{\pi^2}{36}} \]

OpenStudy (anonymous):

well that cncles out the bottom 36 and leaves us with pi^2 so for the top, hold on...

myininaya (myininaya):

so we can write this as: \[\frac{ \frac{36 \cdot \sqrt{3} \pi }{12} - \cdot \frac{36 \cdot 1}{2}}{\frac{36 \cdot \pi^2}{36}}\]

OpenStudy (anonymous):

I get 3radical3 - 1/2

myininaya (myininaya):

unfortunately that isn't correct.

OpenStudy (anonymous):

okay so can you work this out for me? I see what your doing but im probably making tiny mistakes along the way

myininaya (myininaya):

\[\frac{\frac{\cancel{12} \cdot 3 \sqrt{3} \pi }{ \cancel{12}} -\frac{\cancel{2} \cdot 18 }{\cancel{2}}}{\frac{\cancel{36} \pi^2}{\cancel{36}}}\]

myininaya (myininaya):

the oh reason we multiplied by 36 on top and bottom is so we could get rid of mini-fractions and we see here that it worked.

OpenStudy (anonymous):

so we get 3radical3pi-18/pi^2

OpenStudy (anonymous):

Ive gotta get going though, we have been at this for 45 minutes. Thank you for your help

myininaya (myininaya):

Well if you type it in somewhere make sure you write (3sqrt(3)pi-18)/pi^2 Like put parenthesis around the thing you want to be the numerator so people don't think that the numerator is just 18 for that one part of the expression.

myininaya (myininaya):

\[\frac{3 \sqrt{3} \pi -18}{\pi^2} \]